# Density Estimation via Measure Transport: Outlook for Applications in the Biological Sciences

### v2

@@ 6 Discussion
Using such time-dependent gene expression data, one can compute then a sequence of triangular transport maps $\{\mathcal{T}_{k}(t_{l})\}_{l=1}^{n}$, and use them as validation points for a stochastic model describing the dynamics of $\mathcal{T}_{k}(t)$ and thus the correlation time dynamics of the entire gene network after a radiation dose has been applied. Of course, all of the above could be done locally if one restricts the study to a subset of genes, that is, to the time dynamics of a subset of random variables $\mathcal{{Y}}(t)=\left(\mathcal{{Y}}_{1}(t),\mathcal{{Y}}_{2}(t),\dots,\mathcal{{Y}}_{m}(t)\right)\in\mathbb{R}^{m}\subset\mathbb{R}^{d}$ and the corresponding local target probability measure $\nu_{\pi_{k}}(t)$ (or target probability density ${\pi_{k}}(t)$), and the estimating local triangular transport map ${T}_{k}(t)$.
@@
In summary, the work reported in this manuscript shows some of the benefits that measure transport techniques will provide as part of the computational workflow from Section 2, intended to support research in the biological sciences. In domains for which data scarcity is common, such as in the radiation biology case study from Section 5, engagement with domain scientists will aid in identifying types of prior knowledge (in addition to biological pathway knowledge) which can be incorporated into the formulation of the mathematical problems being solved. This should allow to better account for the scarcity of training data. How to incorporate prior knowledge will obviously depend on the type of prior knowledge and problem being addressed but, generally, we mention feature selection for model training as well as incorporation of additional constraints in optimization problems being solved as common approaches.
@@
Immediate tasks to pursue, in collaboration with domain scientists, include generation of synthetic data sets that will allow design of more controlled computational experiments with the goal of gaining a better (quantitative) understanding of the dependence of the methodologies proposed in Section 5.2 on the amount of data available for training. This will aid in carrying out a formal mathematical analysis of the method proposed in Section 5.2.2 as well. Additional (non-synthetic) data sets for validation of the proposed methodologies also will be sought.
@@ Appendix A Appendix: Adaptive Transport Maps Framework
Recall from Section 4.1 of the main text that a transport map $T:\mathbb{R}^{m}\rightarrow\mathbb{R}^{m}$, such that $\nu_{\pi}=T_{\sharp}\nu_{\rho}$, is triangular if the $i$-th component $T_{i}$ of $T$ only depends on the values $\{x_{j}\}_{j=1}^{i}$ of the first $i$ random variables $\{\mathcal{{X}}_{j}\}_{j=1}^{i}$:
@@
Conversely, and denoting the map components of the inverse $T^{-1}:\mathbb{R}^{m}\rightarrow\mathbb{R}^{m}$ of $T$ by $S_{i}\equiv(T^{-1})_{i},i=1,\ldots,m$, one has that the $i$-th map component $S_{i}$ only depends on the values $\{y_{j}\}_{j=1}^{i}$ of the first $i$ random variables $\{\mathcal{{Y}}_{j}\}_{j=1}^{i}$:
@@
To ensure positivity of the densities (refer to equations (3) and (4) of the main text), each map component $T_{i}$ in equation (25) must be differentiable with respect to $x_{i}$, and we must have that $\partial_{x_{i}}T_{i}>0$, for every $i=1,\ldots,m$. Similarly, each component $S_{i}$ in equation (30) of the inverse map $T^{-1}$ of $T$ must satisfy said properties with respect to the variables $y_{i}$ [4, 24].
We now briefly describe the method of constructing and computing (the components of) the triangular transport map as it has been implemented in [2]. Let $\gamma(dy)=\frac{1}{\sqrt{2\pi}}e^{-{y^{2}}/{2}}dy$ be the standard Gaussian measure on $\mathbb{R}$. Consider the space $\mathbb{R}^{\mathbb{N}}$ (sometimes denoted $\mathbb{R}^{\infty}$), where =${\mathbb{N}}=\{1,2,\ldots\}$ is the set of natural numbers, and the set $\mathbb{Z}_{+}^{\mathbb{N}}$, where +=$\mathbb{Z}_{+}=\{0,1,2,\ldots\}$ is the set of non-negative integers. For $y=(y_{1},y_{2},\ldots)\in\mathbb{R}^{\mathbb{N}}$, the canonical projections $P^{(i)}:\mathbb{R}^{\mathbb{N}}\rightarrow\mathbb{R}^{i}$ are defined as $P^{(i)}(y)=(y_{1},y_{2},\ldots,y_{i})$. Similarly, if $\alpha=(\alpha_{1},\alpha_{2},\ldots)\in\mathbb{Z}_{+}^{\mathbb{N}}$, and using the same notation for the canonical projection $P^{(i)}:\mathbb{Z}_{+}^{\mathbb{N}}\rightarrow\mathbb{Z}_{+}^{i}$, we have $P^{(i)}(\alpha)=(\alpha_{1},\alpha_{2},\ldots,\alpha_{i})\equiv\alpha^{(i)}$. Let $\gamma^{\mathbb{N}}=\gamma\times\gamma\times\cdots$ be the infinite product Gaussian measure defined on the Borel $\sigma$-algebra $\mathcal{B}(\mathbb{R}^{\mathbb{N}})$. Then we obtain $P^{(i)}_{\sharp}\left(\gamma^{\mathbb{N}}\right)=\gamma^{i}=\underbrace{\gamma\times\gamma\times\cdots\times\gamma}_{i\,\mathrm{times}}$ to be the $i$-variate Gaussian product measure on the Borel $\sigma$-algebra $\mathcal{B}(\mathbb{R}^{i})$. Let $L^{2}(\mathbb{R}^{i},\gamma^{i})$ be the space of square integrable real valued functions on $\mathbb{R}^{i}$ with respect to $\gamma^{i}$. For all $i\in\mathbb{N}$ consider the Sobolev spaces
@@
Then the set $\{h_{n}\}_{n\geq 0}$ forms an orthonormal basis on the Hilbert space $L^{2}(\mathbb{R},\gamma)$. Moreover, $\{h_{n}\}_{n\geq 0}$ forms an orthogonal basis on the Hilbert space $H^{1}(\mathbb{R},\gamma)$ as well, with its inner product
@@
for every $f,g\in H^{1}(\mathbb{R},\gamma)$. More generally, the space of polynomials is dense in $L^{2}(\mathbb{R},\gamma)$ and this implies that the space of polynomials on $\mathbb{R}^{i}$ is dense in $L^{2}(\mathbb{R}^{i},\gamma^{i})$, for every $i\in\mathbb{N}$. Therefore, with the remark above in mind, we take as an orthonormal basis for the Hilbert space $V_{i}(\mathbb{R}^{i},\gamma^{i})\subset L^{2}(\mathbb{R}^{i},\gamma^{i})$ the set of polynomials
@@
where the polynomials $\prod_{s=1}^{i-1}h_{\alpha_{s}}(y_{s})$, $(\alpha_{1},\alpha_{2},\ldots,\alpha_{i-1})\in\mathbb{Z}_{+}^{i-1}$, form an orthonormal basis of $L^{2}(\mathbb{R}^{i-1},\gamma^{i-1})$ and ${\tilde{h}}_{\alpha_{i}}(y_{i})=h_{\alpha_{i}}(y_{i})/\sqrt{\alpha_{i}+1}$, $\alpha_{i}\in\mathbb{Z}_{+}$, form an orthonormal basis of $H^{1}(\mathbb{R},\gamma)$ (the latter normalization follows from the property $h_{n+1}^{\prime}=\sqrt{n+1}\,h_{n}$ of Hermite polynomials). Finally, let us consider the space
$H^{\infty}(\mathbb{R},\gamma)=\left\{f\in L^{2}(\mathbb{R},\gamma)\mid\forall{s}:s\in\mathbb{N},f^{(s)}\in L^{2}(\mathbb{R},\gamma)\right\}$ (38)
of infinitely differentiable functions on $\mathbb{R}$. Using the definition (35) and integration by parts one can easily show that $\langle f,h_{n}\rangle_{L^{2}(\mathbb{R},\gamma)}={1}/{\sqrt{n!}}\langle f^{(n)},1\rangle_{L^{2}(\mathbb{R},\gamma)}$ and therefore every function $f\in H^{\infty}(\mathbb{R},\gamma)$ can be represented as
$f(y)=\sum_{n=0}^{\infty}\frac{1}{\sqrt{n!}}\left\langle f^{(n)},1\right\rangle_{L^{2}(\mathbb{R},\gamma)}h_{n}(y).$ (39)
This result has a straightforward generalization in arbitrary dimension $i\geq 1$. Namely, every infinitely differentiable function $f\in H^{\infty}(\mathbb{R}^{i},\gamma^{i})$, where
$H^{\infty}(\mathbb{R}^{i},\gamma^{i})=\left\{f\in L^{2}(\mathbb{R}^{i},\gamma^{i})\,\mid\,\alpha^{(i)}\in\mathbb{Z}_{+}^{i}\setminus\{\mathbf{0}\},\frac{\partial^{|\alpha^{(i)}|}f}{\partial y^{\alpha_{1}}\cdots\partial y^{\alpha_{i}}}\in L^{2}(\mathbb{R}^{i},\gamma^{i})\right\},$ (40)
can be represented as
$f(y)=\sum_{|\alpha^{(i)}|=0}^{\infty}\frac{1}{\sqrt{\alpha^{(i)}!}}\left\langle\frac{\partial^{|\alpha^{(i)}|}f}{\partial y^{\alpha_{1}}\cdots\partial y^{\alpha_{i}}},1\right\rangle_{L^{2}(\mathbb{R}^{i},\gamma^{i})}\mathcal{H}_{\alpha^{(i)}}(y),$ (41)
where $|\alpha^{(i)}|=\alpha_{1}+\cdots+\alpha_{i}$, $\alpha^{(i)}!=\alpha_{1}!\cdots\alpha_{i}!$, and the polynomials $\mathcal{H}_{\alpha^{(i)}}(y)=h_{\alpha_{1}}(y_{1})\cdots h_{\alpha_{i}}(y_{i})$, $\alpha^{(i)}\in\mathbb{Z}_{+}^{i}$, form an orthonormal basis for $L^{2}(\mathbb{R}^{i},\gamma^{i})$.
The map components $\{S_{i}\}_{i=1}^{m}$ in equation (30) are sought via rectification operators $\{\mathcal{R}_{i}\}_{i=1}^{m}$ of the form
[…]
that take sufficiently smooth non-monotone functions $f_{i}:\mathbb{R}^{i}\rightarrow\mathbb{R}$, $i=1,\ldots,m$, and transform them into monotone increasing functions $S_{i}$ of their last argument $y_{i}$ (see (30)). In equation (42) the function $g:\mathbb{R}\rightarrow\mathbb{R}^{+}$ is bijective and positive-valued. As discussed in [2], the choice of the function $g$ in equation (42) has significant impact on the properties of the optimization problem (19) (refer to Section 4.1 of the main text) solved in order to compute, or learn, the transport map. The soft-plus function $g(z)=\log(1+\exp(z))$, whose inverse is $g^{-1}(z)=\log(\exp(z)-1)$, is one appropriate choice.
For each $i=1,\ldots,m$, the function $f_{i}$ in equation (42) is represented as an $M_{i}$-term linear combination of basis functions from the set (37) (a finite $M_{i}$-dimensional approximation)
@@
  • $c_{\alpha^{(i)}}\in\mathbb{R}$ are real-valued coefficients (coordinates with respect to the basis), and
@@
Let us now give an outline of the adaptive transport map algorithm applied in our computations. The sets (of active indices which identify the sets of active variables – refer to Figure 3 of the main text) $\Lambda^{(i)}$ in equation (43) are constructed by a greedy algorithm inspired by the work from [26]. Namely, starting with a set $\Lambda^{(i)}_{0}=\emptyset$ one constructs a sequence $\{\Lambda^{(i)}_{t}\}_{t\geq 0}$ of downward closed sets. A set $\Lambda^{(i)}_{t}\subset\mathbb{Z}_{+}^{i}$ is downward closed if $\alpha^{(i)}\in\Lambda^{(i)}_{t}$ and $\alpha^{\prime(i)}\leq\alpha^{(i)}\Longrightarrow\alpha^{\prime(i)}\in\Lambda^{(i)}_{t}$, where $\alpha^{\prime(i)}\leq\alpha^{(i)}$ means $\alpha^{\prime}_{s}\leq\alpha_{s}$, for $1\leq s\leq i$. The reduced margin $\mathcal{RM}(\Lambda^{(i)}_{t})$ of $\Lambda^{(i)}_{t}$ is defined as
@@
Cross-validation [14] is employed to determine the maximal cardinality $M_{i}^{\mathrm{max}}$ of $\Lambda^{(i)}_{t}$ (that is, the step $t=M_{i}^{\mathrm{max}}$ at which to stop adaptation), without overfitting the data. This results in an optimal value $M_{i}=\mathrm{Card}(\Lambda^{(i)})\leq M_{i}^{\mathrm{max}}$ and an approximation $f_{i}$ which is thus adapted to the size of the training sample. The full details are provided in [2]. The greedy, incremental search for values $M_{i}=\mathrm{Card}(\Lambda^{(i)})\leq M_{i}^{\mathrm{max}}$ of the cardinalities of the sets $\Lambda^{(i)}$ leads to numerical efficiency in the algorithm (as the number of unknowns to determine in equation (43) grows quickly with increasing $i$ and $\mathrm{Card}(\Lambda^{(i)})$), as well as numerical stability, particularly when the amount of training data is “small”.

### v4

@@ 6 Discussion
Using such time-dependent gene expression data, one can learn then a sequence of triangular transport maps $\{\mathcal{T}_{k}(t_{l})\}_{l=1}^{n}$, and use them as validation points for a stochastic model describing the dynamics of $\mathcal{T}_{k}(t)$ and thus the correlation time dynamics of the entire gene network after a radiation dose has been applied. Of course, all of the above could be done locally if one restricts the study to a subset of genes, that is, to the time dynamics of a subset of random variables $\mathcal{{Y}}(t)=\left(\mathcal{{Y}}_{1}(t),\mathcal{{Y}}_{2}(t),\dots,\mathcal{{Y}}_{m}(t)\right)\in\mathbb{R}^{m}\subset\mathbb{R}^{d}$ and the corresponding local target probability measure $\nu_{\psi_{k}}(t)$ (or target probability density ${\psi_{k}}(t)$), and the estimating local triangular transport map ${T}_{k}(t)$.
@@
In summary, the work reported in this manuscript shows some of the benefits that measure transport techniques will provide as part of the computational workflow from Section 2, intended to support research in the biological sciences. In domains for which it is common to have limited number of samples available for model training, such as in the radiation biology case study from Section 5, engagement with domain scientists will aid in identifying types of prior knowledge (in addition to biological pathway knowledge) which can be incorporated into the formulation of the mathematical problems being solved. How to incorporate prior knowledge will obviously depend on the type of prior knowledge and problem being addressed but, generally, we mention as common approaches feature selection for model training as well as incorporation of additional constraints in optimization problems being solved.
@@
Immediate tasks to pursue, in collaboration with domain scientists, include generation of synthetic data sets that will allow design of more controlled computational experiments with the goal of gaining a better (quantitative) understanding of the dependence of the methodologies proposed in Section 5.2 on the number of samples available for training. This will aid in carrying out a formal mathematical analysis of the method proposed in Section 5.2.2 as well. Additional (non-synthetic) data sets for validation of the proposed methodologies also will be sought.
@@ Acknowledgements
Statements and Declarations
Funding: This work was supported by the U.S. Department of Energy, Office of Science, RadBio program under Award KP1601011/FWP CC121.
Financial interests: The authors declare they have no financial interests.
Conflict of interests: The authors declare they have no potential conflict of interests.
Data Availability: The data sets [20, 21] used in this research are publicly available.
@@ Appendix A Appendix: Adaptive Transport Maps Framework
Recall from Section 4.1 of the main text that a transport map $T\!:\!\mathbb{R}^{m}\!\rightarrow\!\mathbb{R}^{m}$, such that the measure transport equation $\nu_{\psi}=T_{\sharp}\nu_{\rho}$ holds, is triangular if the $i$-th component $T_{i}$ of $T$ depends only on the values $\{x_{j}\}_{j=1}^{i}$ of the first $i$ random variables $\{\mathcal{{X}}_{j}\}_{j=1}^{i}$:
@@
Conversely, and denoting the map components of the inverse $T^{-1}\!:\!\mathbb{R}^{m}\!\rightarrow\!\mathbb{R}^{m}$ of $T$ by $S_{i}\equiv(T^{-1})_{i},i=1,\ldots,m$, one has that the $i$-th map component $S_{i}$ depends only on the values $\{y_{j}\}_{j=1}^{i}$ of the first $i$ random variables $\{\mathcal{{Y}}_{j}\}_{j=1}^{i}$:
@@
Recall also that, to ensure positivity of the densities (refer to equations (6) and (7) of the main text), each map component $T_{i}$ in equation (33) must be differentiable with respect to $x_{i}$, and we must have that $\partial_{x_{i}}T_{i}>0$, for every $i=1,\ldots,m$. Similarly, each component $S_{i}$ in equation (38) of the inverse map $T^{-1}$ of $T$ must satisfy said properties with respect to the variables $y_{i}$ [4, 24].
We now move on to describe briefly the method of constructing and computing (the components of) a triangular transport map as it has been implemented in [3]. First, we set the stage to define the space(s) of functions whose elements can be used as building blocks for the construction of the components of the transport map and whose bases will be used to construct approximations to the transport map components (via equation (46), later in this section).
To begin, let $\gamma(\mathrm{d}{{y}})=\frac{1}{\sqrt{2\pi}}e^{-{y^{2}}/{2}}\mathrm{d}{{y}}$ be the standard Gaussian measure on $\mathbb{R}$. Consider the space $\mathbb{R}^{\mathbb{N}}$ (sometimes denoted $\mathbb{R}^{\infty}$), where =${\mathbb{N}}=\{1,2,\ldots\}$ is the set of natural numbers, and the set $\mathbb{Z}_{+}^{\mathbb{N}}$, where +=$\mathbb{Z}_{+}=\{0,1,2,\ldots\}$ is the set of non-negative integers. For $y=(y_{1},y_{2},\ldots)\in\mathbb{R}^{\mathbb{N}}$, the canonical projections $P^{(i)}:\mathbb{R}^{\mathbb{N}}\rightarrow\mathbb{R}^{i}$ are defined as $P^{(i)}(y)=(y_{1},y_{2},\ldots,y_{i})$. Similarly, if $\alpha=(\alpha_{1},\alpha_{2},\ldots)\in\mathbb{Z}_{+}^{\mathbb{N}}$, and using the same notation for the canonical projection $P^{(i)}:\mathbb{Z}_{+}^{\mathbb{N}}\rightarrow\mathbb{Z}_{+}^{i}$, we have $P^{(i)}(\alpha)=(\alpha_{1},\alpha_{2},\ldots,\alpha_{i})\equiv\alpha^{(i)}$. Let $\gamma^{\mathbb{N}}=\gamma\times\gamma\times\cdots$ be the infinite product Gaussian measure defined on the Borel $\sigma$-algebra $\mathcal{B}(\mathbb{R}^{\mathbb{N}})$. Then we obtain $P^{(i)}_{\sharp}\left(\gamma^{\mathbb{N}}\right)=\gamma^{i}=\underbrace{\gamma\times\gamma\times\cdots\times\gamma}_{i\,\mathrm{times}}$ to be the $i$-variate Gaussian product measure on the Borel $\sigma$-algebra $\mathcal{B}(\mathbb{R}^{i})$. Let $L^{2}(\mathbb{R}^{i},\gamma^{i})$ be the space of square integrable real valued functions on $\mathbb{R}^{i}$ with respect to $\gamma^{i}$. For all $i\in\mathbb{N}$ consider the Sobolev spaces
@@
The set $\{h_{n}\}_{n\geq 0}$ forms an orthonormal basis on the Hilbert space $L^{2}(\mathbb{R},\gamma)$. Moreover, $\{h_{n}\}_{n\geq 0}$ forms an orthogonal basis on the Hilbert space $H^{1}(\mathbb{R},\gamma)$ as well, with its inner product
@@
for every $f,g\in H^{1}(\mathbb{R},\gamma)$. More generally, the space of polynomials is dense in $L^{2}(\mathbb{R},\gamma)$ and this implies that the space of polynomials on $\mathbb{R}^{i}$ is dense in $L^{2}(\mathbb{R}^{i},\gamma^{i})$, for every $i\in\mathbb{N}$. Therefore, with the latter remark in mind, we take as an orthonormal basis for the Hilbert space $V_{i}(\mathbb{R}^{i},\gamma^{i})\subset L^{2}(\mathbb{R}^{i},\gamma^{i})$ from (41) the set of polynomials
@@
where the polynomials $\prod_{s=1}^{i-1}h_{\alpha_{s}}(y_{s})$, $(\alpha_{1},\alpha_{2},\ldots,\alpha_{i-1})\in\mathbb{Z}_{+}^{i-1}$, form an orthonormal basis of $L^{2}(\mathbb{R}^{i-1},\gamma^{i-1})$ and ${\tilde{h}}_{\alpha_{i}}(y_{i})=h_{\alpha_{i}}(y_{i})/\sqrt{\alpha_{i}+1}$, $\alpha_{i}\in\mathbb{Z}_{+}$, form an orthonormal basis of $H^{1}(\mathbb{R},\gamma)$ (the latter normalization follows from the property $h_{n+1}^{\prime}=\sqrt{n+1}\,h_{n}$ of Hermite polynomials). Then, for each $i=1,\ldots,m$, the elements of the space $V_{i}(\mathbb{R}^{i},\gamma^{i})$ with its basis (45) can be used as building blocks for constructing the $i$-th component $S_{i}$ of the triangular transport map as well as its approximations.
With the mathematical formalism outlined above, we can now describe the functional approximations used to represent the triangular transport maps. The map components $\{S_{i}\}_{i=1}^{m}$ in equation (38) are sought via rectification operators $\{\mathcal{R}_{i}\}_{i=1}^{m}$ of the form
$S_{i}(y_{1},\ldots,y_{i})\,=\,\mathcal{R}_{i}(f_{i})(y_{1},\ldots,y_{i})\,=\,f_{i}(y_{1},\ldots,y_{i-1},0)+\int_{0}^{y_{i}}g\left(\partial_{t}f_{i}(y_{1},\ldots,y_{i-1},t)\right)\,\mathrm{d}{{t}},$ (46)
that take sufficiently smooth (e.g., $f_{i}\in V_{i}(\mathbb{R}^{i},\gamma^{i})$ – see (41)) non-monotone functions $f_{i}:\mathbb{R}^{i}\rightarrow\mathbb{R}$, $i=1,\ldots,m$, and transform them into monotone increasing functions $S_{i}$ of their last argument $y_{i}$ (see (38)). In equation (46) the function $g:\mathbb{R}\rightarrow\mathbb{R}^{+}$ is bijective and positive-valued. This guarantees that the positivity of the densities (equations (6) and (7) of the main text) is not violated, a condition that must be satisfied. As discussed in [3], the choice of the function $g$ in equation (46) has significant impact on the properties of the optimization problem (24) (refer to Section 4.1 of the main text) solved in order to compute, or learn, the transport map. The soft-plus function $g(z)=\log(1+\exp(z))$, whose inverse is $g^{-1}(z)=\log(\exp(z)-1)$, is one appropriate choice.
For each $i=1,\ldots,m$, the function $f_{i}$ in equation (46) is represented as an $M_{i}$-term linear combination of basis functions from the set (45) (a finite $M_{i}$-dimensional approximation),
@@
  • $c_{\alpha^{(i)}}\in\mathbb{R}$ are real-valued coefficients (coordinates with respect to the basis), which are to be learned from data, and
@@
Finally, let us give an outline of the adaptive transport map algorithm applied in our computations to learn the transport maps. The sets (of active indices which identify the sets of active variables – refer to Figure 3 of the main text) $\Lambda^{(i)}$ in equation (47) are constructed by a greedy algorithm inspired by the work from [26]. Namely, starting with a set $\Lambda^{(i)}_{0}=\emptyset$ one constructs a sequence $\{\Lambda^{(i)}_{t}\}_{t\geq 0}$ of downward closed sets. A set $\Lambda^{(i)}_{t}\subset\mathbb{Z}_{+}^{i}$ is downward closed if $\alpha^{(i)}\in\Lambda^{(i)}_{t}$ and $\alpha^{\prime(i)}\leq\alpha^{(i)}\Longrightarrow\alpha^{\prime(i)}\in\Lambda^{(i)}_{t}$, where $\alpha^{\prime(i)}\leq\alpha^{(i)}$ means $\alpha^{\prime}_{s}\leq\alpha_{s}$, for $1\leq s\leq i$. The reduced margin $\mathcal{RM}(\Lambda^{(i)}_{t})$ of $\Lambda^{(i)}_{t}$ is defined as
@@
Cross-validation [14] is employed to determine the maximal cardinalities $M_{i}^{\mathrm{max}}$ of each set $\Lambda^{(i)}_{t}$ (that is, the step $t=M_{i}^{\mathrm{max}}$ at which to stop adaptation, for each $i=1,\ldots,m$), without overfitting the training data. This results in values $M_{i}=\mathrm{Card}(\Lambda^{(i)})\leq M_{i}^{\mathrm{max}}$ and approximations $f_{i}$ which are optimally adapted to the size of the training data set. The full details are provided in [3]. The greedy, incremental search for values $M_{i}=\mathrm{Card}(\Lambda^{(i)})\leq M_{i}^{\mathrm{max}}$ of the cardinalities of the sets $\Lambda^{(i)}$ leads to numerical efficiency in the algorithm (as the number of unknowns to determine in equation (47) grows quickly with increasing $i$ and $\mathrm{Card}(\Lambda^{(i)})$), as well as numerical stability in the computations, including when training with a limited number of data samples.